OpenStudy (anonymous):

find the limit as x->0 of (1-cos3h)/((cos^2)5h-1)

OpenStudy (anonymous):

just asking, did you double check the numbers?

OpenStudy (anonymous):

yeah, just to be clear: find the limit as x -> 0 \[\frac{ 1-\cos3h }{ \cos ^{2}5h - 1}\]

OpenStudy (isaiah.feynman):

And there is no x in that expression!

OpenStudy (anonymous):

lol

OpenStudy (isaiah.feynman):

Perhaps you mean h. @j88clark

OpenStudy (anonymous):

yeah

OpenStudy (isaiah.feynman):

The limit is a 0/0 type which is an indeterminate form. So use the L'hospital's rule.

OpenStudy (anonymous):

were supposed to be transforming it with identities

OpenStudy (anonymous):

I think, @Isaiah.Feynman is correct, I cannot think of another way as well

OpenStudy (isaiah.feynman):

the denominator can be re written as \[-\sin^{2}5h\]

OpenStudy (isaiah.feynman):

@j88clark

OpenStudy (isaiah.feynman):